# Expert-elicitation method for non-parametric joint priors using normalizing flows

### v1

@@ 4.2 Simulation Study 1: Binomial Likelihood with Independent Normal Priors
We introduce the general idea of our method using a Binomial model with a logit link (sigmoid response function) and one continuous predictor $x=1,\ldots,50$, scaled by its standard deviation. In the following we refer to this model as M1:
@@
As elicited statistics, we select two observations from the continuous predictor $x$, denoted by $x_{0}$ and $x_{1}$, which correspond to the $25\%$ and $75\%$ quantiles of $x_{i}$. We select specific quantiles rather than randomly sampling two observations from the predictor to avoid the observations being too close together, which would reduce informativeness. From the two selected observations, we compute five quantiles from the prior predictive distribution to represent the elicited statistics (see Section 4.1 for details). To obtain the “expert”-elicited statistics, we define a true prior that represents the ground truth and simulate from the generative model in forward mode, computing the corresponding true-elicited statistics. The true joint prior is defined by independent normal distributions for each model parameter: $\beta_{0}\sim\text{Normal}(0.1,0.1)$ and $\beta_{1}\sim\text{Normal}(-0.1,0.3)$.
To assess the informativeness of the selected set of elicited statistics, we conduct a sensitivity analysis, with the results shown in Figure 3. In each row, a single hyperparameter is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The columns present the two elicited statistics, specifically the five quantiles (in shades of blue and green) for $y\mid x_{0}$ and $y\mid x_{1}$. The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$. The results of the sensitivity analysis indicate that the selected elicited statistics provide sufficient informativeness for the model hyperparameters.
Figure 3: Binomial model—Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid x_{0}$ and $y\mid x_{1}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
@@
We ran the training algorithm 30 times with different random seeds and conducted an initial convergence check by inspecting the slope of the last 100 epochs for each replication. In Figure 4, the upper row shows the absolute slope (scaled by a factor of 100) for each replication. In the lower row, the learning trajectory of the total loss is shown for the best-performing (leftmost) and the four worst-performing replications. We observe a near-zero slope for the best model, indicating convergence, while the negative slopes for the worst-performing seeds suggest that the learning algorithm continues to make progress. However, a visual inspection of the learning trajectories for the four worst-performing replications, exemplified in Figure 5 for seed 11, indicates satisfactory convergence for all quantities of interest. This includes the total loss, the individual loss components (upper row), as well as the means and standard deviations of the marginal priors (lower row). Therefore, although there appears to be minor ongoing learning progress, the results will only change very minimally. Overall, we conclude that the training process was successful.
@@
Figure 4: Binomial model—Preliminary convergence check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 28, while the worst-performing models, indicated by negative slopes, are seeds 3, 14, 15, and 11. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
@@
Figure 5: Binomial model—Visual convergence check for seed=11: The plots display the updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
@@
The corresponding final learned elicited statistics for each replication are shown in the first two plots in Figure 6. Each true quantile (x-axis) is plotted against its corresponding learned quantile (y-axis) for each replication. Points that align perfectly with the diagonal dashed line represent a perfect match between true and learned quantiles. Points above the diagonal indicate that the learned quantiles are higher than the true ones (and vice versa for lower points). Additionally, in the rightmost plot in Figure 6, we present the true (red) and learned (orange) correlations between the marginal priors. Consistent with successful convergence, the elicited statistics, including the correlations, are accurately learned across all replications.
@@
Figure 6: Binomial model—Learned elicited statistics for all 30 replications: The two leftmost plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the two target quantities, $y\mid x_{0}$ and $y\mid x_{1}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (orange) vs. true (red) correlations. Given that we assume independence between model parameters in this scenario, the true correlation is zero. Consistent with convergence, all learned elicited statistics closely match the ground truth.
@@
Finally, we can examine the prior distribution learned by each of the 30 replications. Figure 7b displays the learned marginal priors for each replication in grey. Since all priors closely align with the ground truth (indicated by the dotted black line), the different seeds are visually indistinguishable. These results suggest that the set of elicited statistics provides sufficient information to practically identify model M1. This conclusion is further supported by the weights used for model averaging, depicted in Figure 7a. The weights represent the relative performance of each replication based on their total loss (see Section 2.5 for details). All replications perform equally well, resulting in a uniform weighting scheme. The corresponding model average is depicted via a dashed black line in Figure 7b. In this example, all learned marginal priors, the ground truth, and the model average coincide, showcasing an ideal scenario.
Figure 7: Binomial model—Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. In this scenario, all replications perform equally well, resulting in uniform weights. (b) The learned marginal priors for each replication are shown in grey. The dotted black line represents the ground truth and the dashed, black line indicates the model average. In this case, the elicited statistics are sufficiently informative to practically identify the generative model, leading the learned priors, the ground truth, and the model average to coincide.
@@ 4.3 Simulation Studies 2: Normal Likelihood
In the remaining simulation studies, we aim to evaluate the method’s performance using a model with increased complexity, specifically a normal regression model with standard deviation $\sigma$ and mean $\mu_{i}$, where $\mu_{i}$ depends on a three-level grouping variable encoded via the dummy variables $x_{1}$ and $x_{2}$:
@@
  • Scenario 2 (M3): This setup is identical to M2 but introduces skewed normal priors for $\beta_{1}$ and $\beta_{2}$, where $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$ [^1] For the skewed normal distribution, we use the implementation from TensorFlow Probability [1], which features the two-piece normal distribution [24]. This distribution is parameterized by location, scale, and shape parameters. The shape parameter controls the skewness, with values above one resulting in a right-skewed distribution. and $\beta_{2}\sim\text{SkewNormal}(2.5,0.8,4)$.
  • Scenario 3 (M4): This scenario introduces a dependency structure among the model coefficients $\beta$ using a correlated multivariate normal distribution, with marginal distributions identical to those in M2. All other aspects remain the same as in M2. The correlated prior is specified as:
@@ 4.3.1 Scenario 1: Independent Normal Priors
We begin with a brief assessment of the selected elicited statistics, with sensitivity analysis results detailed in Appendix A.2. While the intercept coefficient, $\beta_{0}$, is well-informed by all three target quantities related to the predictive distributions of the groups, the slope coefficients, $\beta_{1}$ and $\beta_{2}$, are each informed by only one target quantity, specifically those corresponding to group 2 and group 3, respectively. Consequently, we anticipate greater difficulty in accurately learning these coefficients. Additionally, we included $R^{2}$ among the target quantities to aid the learning algorithm in distinguishing parameter uncertainty from data uncertainty. However, given the limited data available, we expect considerable variability in the learned model variances.
@@
To verify successful convergence, we first examined the absolute slopes across all 30 replications, followed by a detailed visual inspection of the learning trajectories for the loss components and additional quantities of interest in the replications with the highest absolute slopes (see Appendix A.2 for results). Based on this analysis, we conclude that the learning process was successful. The corresponding learned model-implied statistics are shown in Figure 8 and demonstrate a strong match between the learned and true elicited statistics.
Figure 8: Normal model / Scenario 1 (M2) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero. All learned elicited statistics closely match the ground truth.
@@
The corresponding learned marginal priors of each replication are depicted in Figure 9. Figure 9a displays the distribution of the computed weights used for model averaging. The replication with the highest weight is highlighted in green, while the replication with the lowest weight is highlighted in orange. In Figure 9b, the learned marginal priors of each replication are depicted in grey. The dotted black line refers to the ground truth and the dashed black line to the model average. Additionally, the marginal priors for the best- and worst-performing replications are highlighted in their respective colors. The results suggest that the elicited statistics provide sufficient information to accurately learn the marginal prior of $\beta_{0}$. In contrast, the remaining learned marginal priors display substantial variation. Notably, for $\sigma$, most replications yield a marginal prior concentrated around the mean of a Gamma(5, 2) distribution. This indicates that, while the elicited statistics capture some information about the average level of data variation, they convey limited information about its uncertainty. However, despite the variability in the learned marginal priors, there is no degenerate or otherwise highly outlying distribution whose plausibility we would exclude a priori.
Figure 9: Normal model / Scenario 1 (M2) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line indicating the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors. Although the elicited statistics provide sufficient information to accurately learn the prior for $\beta_{0}$, the substantial variation in the learned priors for the other model parameters indicates a certain lack of model identification.
@@ 4.3.2 Scenario 2: Skewed Normal Priors
This scenario differs from the previous one only in the ground truth, where we introduce a skew-normal marginal prior for the two slope coefficients, $\beta_{1}$ and $\beta_{2}$. However, since we use the same set of ”elicited” statistics, the sensitivity analysis (see Appendix A.3) already suggests that there is likely insufficient information to disentangle the variance and skewness components for these two slope coefficients; therefore, these components are most likely not identified.
Fitting Stage
[…]
Figure 10: Normal model / Scenario 2 (M3) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero. All learned elicited statistics closely match the ground truth, except for $R^{2}$, where the learned quantiles tend to be overestimated, implying an underestimation of $\sigma$ (see text for details).
Evaluation Stage
Indeed, when examining the learned marginal priors in Figure 10b, we see that for $\sigma$, there is one replication that places a significant amount of prior mass on values near zero. This replication corresponds to the seed with the smallest weight as depicted in Figure 10a and is therefore highlighted in orange. The next step in this case would be to consult with the domain expert on the plausibility of the learned marginal priors, particularly considering whether the marginal prior for $\sigma$, with values close to zero, is meaningful in the context of the study. Since $\sigma$ shares the same scale as the outcome variable $y$, it allows for an assessment of how much or how little noise in the data is considered realistic. This situation illustrates how, even if a domain expert may not have been able to express precise expectations for $\sigma$ initially, the resulting distributions can enable them to retrospectively rule out certain cases they find unrealistic.
Figure 11: Normal model / Scenario 2 (M3) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors. Substantial variation in the learned priors for all model parameters suggests a lack of model identification. However, no degenerate distribution is learned. For $\sigma$, the learned priors exhibit some degree of multi-modality, with one replication in particular placing a significant amount of probability mass on small values of $\sigma$.

### v2

@@ 4.2 Simulation Study 1: Binomial Likelihood with Independent Normal Priors
We introduce the general idea of our method using a Binomial model with a logit link (sigmoid response function) and a single continuous predictor, $x$. The predictor was generated by first constructing a sequence from 1 to 50, which was then scaled by its standard deviation. In the following we refer to this model as M1:
@@
As elicited statistics, we select prior predictions for two values of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$. The two values of $x$ correspond to the $25\%$ and $75\%$ quantiles of the predictor. We select specific quantiles rather than randomly sampling two observations from the predictor to avoid the observations being too close together, which would reduce informativeness. From the distribution of the two selected observations, we compute five quantiles which represent the elicited statistics (see Section 4.1 for details). Note that $y\mid x_{0}$ and $y\mid x_{1}$ each have an associated distribution, resulting from simulating $y^{(s)}$ for $S$ samples drawn from the joint prior and conditioning on the respective $x$ value. The distribution over $y\mid x$ encodes the parameter uncertainty.
To obtain the “expert”-elicited statistics, we define a true prior that represents the ground truth and simulate from the generative model in forward mode, computing the corresponding true-elicited statistics. The true joint prior is defined by independent normal distributions for each model parameter: $\beta_{0}\sim\text{Normal}(0.1,0.1)$ and $\beta_{1}\sim\text{Normal}(-0.1,0.3)$.
To assess the informativeness of the selected set of elicited statistics, we conduct a sensitivity analysis, with the results shown in Figure 2. In each row, a single hyperparameter is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The columns present the two elicited statistics, specifically the five quantiles (in shades of blue and green) for $y\mid x_{0}$ and $y\mid x_{1}$. The upper two rows show results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$. The results of the sensitivity analysis indicate that the selected elicited statistics provide sufficient informativeness for the model hyperparameters.
Figure 2: Binomial model—Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid x_{0}$ and $y\mid x_{1}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The upper two rows indicate results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
@@
We ran the training algorithm 30 times with different random seeds and conducted an initial convergence check by inspecting the slope of the last 100 epochs for each replication. In Figure 3, the upper row shows the absolute slope (scaled by a factor of 100) for each replication. In the lower row, the learning trajectory of the total loss is shown for the best-performing (leftmost) and the four worst-performing replications. We observe a near-zero slope for the best model, indicating convergence, while the negative slopes for the worst-performing seeds suggest that the learning algorithm continues to make progress.
However, a visual inspection of the learning trajectories of the total loss and individual loss components (Figure 5) and the means and standard deviations of the marginal priors (Figure 5), indicates satisfactory convergence. Therefore, although the slope analysis indicates minor ongoing learning progress, no substantial changes in the results are expected if the model is trained for additional epochs. Overall, we conclude that the training process was successful.
@@
Figure 3: Binomial model—Preliminary convergence check for all 30 replications: Upper plot: Absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. Replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. Best-performing model is seed 25, while the worst-performing models are seeds 18, 11, 5, and 0. Lower plot: Loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
@@
Figure 4: Binomial model—Visual convergence check 1: Updating trajectory across epochs for all 30 replications. On the left, trajectory of the total loss and on the right, trajectory of the weighted individual loss components.
Figure 5: Binomial model—Visual convergence check 2: Updating trajectory across epochs for for all 30 replications. Trajectory of the mean (left plot) and standard deviation (right plot) of the marginal priors.
@@
The corresponding final learned elicited statistics for each replication are shown in Figure 6. In the first two plots, each true quantile (x-axis) is plotted against its corresponding learned quantile (y-axis) for each replication. Points that align perfectly with the diagonal dashed line represent a perfect match between true and learned quantiles. Points above the diagonal indicate that the learned quantiles are higher than the true ones (and vice versa for lower points). Accuracy of the learned correlation information is presented in Figure 6. The true (red) correlation matches perfectly with the learned (black) correlations.
@@
Figure 6: Binomial model—Learned elicited statistics for all 30 replications: The two leftmost plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the two target quantities, $y\mid x_{0}$ and $y\mid x_{1}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (black) vs. true (red) correlations. Given that we assume independence between model parameters in this scenario, the true correlation is zero.
@@
Finally, we examine the learned prior distributions. Figure 7 shows the learned marginal priors for each replication, alongside the results from the model averaging approach. The weights used for model averaging are shown in the upper plot of Figure 7 and reflect the relative performance of each replication based on their total loss (see Section 4.1 for details). In this case, all replications perform similarly well, resulting in an almost uniform weighting scheme. The resulting averaged prior is represented by the solid red line in the lower plot, shown together with the learned priors from each replication.
Figure 7: Binomial model—Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. In this scenario, all replications perform equally well, resulting in almost uniform weights. Lower plot: The learned marginal priors for each replication and in red the prior average. In this case, the elicited statistics are sufficiently informative to effectively identify the generative model, resulting in only minor variations in the learned priors across replications.
@@ 4.3 Simulation Studies 2: Normal Likelihood
In the remaining simulation studies, we aim to evaluate the method’s performance using a model with slightly increased complexity, specifically a normal regression model with standard deviation $\sigma$ and mean $\mu_{i}$, where $\mu_{i}$ depends on a three-level grouping variable encoded via the dummy variables $x_{1}$ and $x_{2}$:
@@
  • Scenario 2 (M3): Identical to M2 but introduces skewed normal priors for $\beta_{1}$ and $\beta_{2}$, where $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$ [^1] For the skewed normal distribution, we use the implementation from TensorFlow Probability [1], which features the two-piece normal distribution [26]. This distribution is parameterized by location, scale, and shape parameters. The shape parameter controls the skewness, with values above one resulting in a right-skewed distribution. and $\beta_{2}\sim\text{SkewNormal}(2.5,0.8,4)$.
  • Scenario 3 (M4): Identical to M2 but introduces a dependency structure among the model coefficients $\beta$ using a correlated multivariate normal distribution, with marginal distributions identical to those in M2. The correlated prior is specified as:
@@ 4.3.1 Scenario 1: Independent Normal Priors
We begin with a brief assessment of the selected elicited statistics, with sensitivity analysis results detailed in Appendix A.2. While the intercept coefficient, $\beta_{0}$, is well-informed by all three target quantities related to the predictive distributions of the groups, the slope coefficients, $\beta_{1}$ and $\beta_{2}$, are each informed by only one target quantity, specifically those corresponding to group 2 and group 3, respectively. Consequently, we anticipate greater difficulty in accurately learning these coefficients. Additionally, we included $R^{2}$ among the target quantities to aid the learning algorithm in distinguishing parameter uncertainty from data uncertainty.
@@
To verify successful convergence, we first examined the absolute slopes across all 30 replications, followed by a detailed visual inspection of the learning trajectories for the loss components and additional quantities of interest (see Appendix A.2 for results). The analysis indicates that the learning process was successful. The model-implied statistics are shown in Figure 8 and demonstrate a strong match between the learned and true elicited statistics, supporting the assumption of successful convergence.
Figure 8: Normal model / Scenario 1 (M2) — Learned elicited statistics for all 30 replications: First three plots (from the left) show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (black) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero.
@@
The corresponding learned marginal priors of each replication alongside with the model averaging results are depicted in Figure 9. In this scenario, we observe noticeably greater variation in the learned priors across different replications, particularly for $\beta_{1}$ and $\beta_{2}$. This variation is also reflected in the model averaging weights, which now deviate clearly from the uniform distribution observed in Simulation Study 1. These results suggest that while the elicited statistics provide sufficient information to learn the marginal priors of $\beta_{0}$ and $\sigma$, they are insufficient to uniquely identify the priors for $\beta_{1}$ and $\beta_{2}$. It is important to emphasize that the only information available to the method for learning the prior distributions is based on the elicited statistics shown in Figure 8. Given this limited set of information, learning a non-parametric joint prior becomes a highly underconstrained problem, where multiple prior distributions can be consistent with the same set of elicited statistics. However, this should not be interpreted as a limitation of the method itself, but rather as a consequence of the limited information available to the method for learning the joint prior.
Figure 9: Normal model / Scenario 1 (M2) — Learned marginal priors across replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. Lower plot: The learned marginal priors for each replication (in black) and the prior average (in red).
@@ 4.3.2 Scenario 2: Skewed Normal Priors
Scenario 2 differs from the previous one only in the ground truth, where we introduce a skew-normal marginal prior for the two slope coefficients, $\beta_{1}$ and $\beta_{2}$. In this scenario, we examine whether the method can learn this additional skewness property using the same “elicited statistics” as those introduced in the previous case study. The sensitivity analysis (see Appendix A.3) already suggests that there is likely insufficient information to disentangle the variance and skewness components for the two slope coefficients; therefore, we except to observe more variation in the learned prior distributions.
Fitting & Evaluation Stage
The convergence diagnostics are depicted in Appendix A.3 and indicate successful convergence. The corresponding match between the model-implied and true “elicited” statistics is shown in Figure 10, demonstrating overall good performance.
Figure 10: Normal model / Scenario 2 (M3) — Learned elicited statistics for all 30 replications: First four plots from the left depict the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (black) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero.
Figure 11 shows the corresponding learned marginal priors together with the results from prior averaging. The results indicate relatively stable learning of the prior distributions across all replications. As before, the priors for the slope coefficients exhibit higher variation. Most importantly, however, the method is able to learn the introduced skewness.
Figure 11: Normal model / Scenario 2 (M3) — Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. Lower plot: The learned marginal priors for each replication (in black) and the prior average (in red).